# Reinforced Imitative Trajectory Planning for Urban Automated Driving

### v1

@@ Abstract
Reinforcement learning (RL) faces challenges in trajectory planning for urban automated driving due to the poor convergence of RL and the difficulty in designing reward functions. The convergence problem is alleviated by combining RL with supervised learning. However, most existing approaches only reason one step ahead and lack the capability to plan for multiple future steps. Besides, although inverse reinforcement learning holds promise for solving the reward function design issue, existing methods for automated driving impose a linear structure assumption on reward functions, making them difficult to apply to urban automated driving. In light of these challenges, this paper proposes a novel RL-based trajectory planning method that integrates RL with imitation learning to enable multi-step planning. Furthermore, a transformer-based Bayesian reward function is developed, providing effective reward signals for RL in urban scenarios. Moreover, a hybrid-driven trajectory planning framework is proposed to enhance safety and interpretability. The proposed methods were validated on the large-scale real-world urban automated driving nuPlan dataset. The results demonstrated the significant superiority of the proposed methods over the baselines in terms of the closed-loop metrics. The code is available at https://github.com/Zigned/nuplan_zigned.
@@ 1 Introduction
RL has been successfully applied in highway driving tasks, such as lane keeping [8, 9], lane changing [10, 11], and traffic merging [12], where the agents are usually responsible for steering, throttle, and braking control according to vectorized traffic information (e.g., relative positions and relative velocities). RL has also been effectively implemented in urban scenarios. Urban scenarios involve more complex situations than highway scenarios, attributed to intricate road structures and diverse traffic participants with different behavioral patterns. Therefore, the state representation for urban scenarios is typically more complex, leading to an increase in both the parameter scale and structural complexity of neural networks used in RL. Consequently, the gradients computed by RL become insufficient to optimize such deep neural networks [13]. Tackling this issue, RL was combined with supervised learning. For instance, an IL agent was pre-trained to regularize the RL agent training by imposing the Kullback-Leibler divergence between the two agents’ policies in Ref. [14]. The RL agent determined a longitudinal target speed and a target lane according to bird’s-eye-view (BEV) rasterized images, which incorporate map information and relative states between traffic participants. Roach [15] trains an RL agent with privileged BEV semantic segmentation images. Then, the RL agent is employed to collect a dataset, with which a downstream IL agent learns to control the steering, throttle, and braking.
However, most existing RL-based methods [16, 17, 13, 7] for AD only reason one step ahead and are not capable of planning for multiple future steps, which can be necessary to handle complex urban scenarios. Although an RL-based trajectory planning method, trajectory value learning (TRAVL), was proposed in Ref. [7], it has several drawbacks. First, TRAVL selects a trajectory from a finite trajectory set generated by predefined rules, which may reduce trajectory optimality. Second, generating the trajectory set and evaluating every trajectory online increase the computational complexity. Third, the applied scenarios of TRAVL were highway scenarios, while urban scenarios are much more complex.
@@ 2.1 Preliminaries
For high-dimensional problems, such as AD, the value function is usually approximated by a neural network $Q_{\theta}(s,a)$ parameterized by $\theta$. The deep Q-learning algorithm [24] optimizes the approximator $Q_{\theta}(s,a)$ to match a target $q$ output by a secondary frozen target network $Q_{\theta^{\prime}}(s,a)$
@@ 2.2 Framework
As mentioned in Section 1, training a deep neural network-based policy for urban trajectory planning using conventional RL methods is intractable. Therefore, we integrate IL into RL with an actor-critic setting, proposing the reinforced imitative trajectory planning method, as shown in Fig. 1. Specifically, at each time step $t$, the actor takes an action $a_{t}^{0}$ given a state $s_{t}$. Concurrently, the action sampler generates $S$ actions, i.e., $\{a_{t}^{i}\}_{i=1}^{S}$, for the same state $s_{t}$. The actions are defined as trajectories $\{a_{t}^{i}\mid a_{t}^{i}=\{\mathbf{p}_{t+j}^{i}\}_{j=1}^{T_{\text{p}}}\}_{i=0}^{S}$, with spatial location $\mathbf{p}_{t+j}^{i}=(x^{i}_{t+j},y^{i}_{t+j})$ and planning time horizon $T_{\text{p}}$. Afterward, only the action $a_{t}^{0}$ is sent to the environment, which then outputs the reward $r_{t+1}$ and the subsequent state $s_{t+1}$. Next, observing the state $s_{t+1}$, the actor and action sampler output actions $\{a_{t+1}^{i}\}_{i=0}^{S}$, and the process repeats accordingly.
@@
While the actor and the action sampler interact with the environment, the experience $(s_{t},\{a_{t}^{i}\}_{i=0}^{S},r_{t+1},s_{t+1})$ is added to the replay buffer at each time step. Subsequently, a mini-batch of experiences is sampled from the replay buffer to update the critic network. Afterward, for each experience, the action with the highest value estimated by the critic network is selected from $\{a_{t}^{i}\}_{i=0}^{S}$ as the optimal action $a_{t}^{*}$. Finally, the optimal state-action pair $(s_{t},a_{t}^{*})$ is used to update the actor via IL, and the updated actor is employed to interact with the environment at the next time step.
@@ 2.3 Environment
The environment can be built on a real-world urban automated driving dataset. During training, the environment is in non-reactive log-replay mode, where the other traffic participants, such as vehicles, pedestrians, and cyclists, do not react to the behavior of the ego agent and simply act according to the logged data. Upon receiving the action $a_{t}^{0}$ from the actor, the environment propagates the motion of the ego vehicle using an LQR-based trajectory tracker [26] and the kinematic bicycle model.
@@ 2.6 Action Sampler
Ideally, the actor should imitate the optimal action $a^{*}$ that maximizes the critic given a state $s$, i.e., $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$. However, finding the optimal action is our goal in the first place. It is intractable because the action space $\mathcal{A}$ for AD is continuous, high-dimensional, and with kinematic and dynamic constraints. Therefore, TRAVL [7] samples spline-based actions that are physically feasible for the ego agent in the Frenet frame of the road [29]. Inspired by TRAVL, we propose sampling actions based on piece-wise polynomials, allowing more flexible driving maneuvers.
@@
Figure 4: Sampling actions based on piece-wise polynomials
@@ 2.8 Training Objectives
In practice, the MotionFormer is optimized to predict not only the ego agent’s trajectory but also the trajectories of surrounding agents, whose supervisory signals are derived from the dataset. Because the predicted trajectories for surrounding agents are useful for the post-optimization (see Section 4).
@@ 3.1 Model Structure
Bayesian RewardFormer, as shown in Fig. 16, has the same architecture as the actor network’s encoder except for two modifications. First, the temporal attention block is removed because Bayesian RewardFormer takes as input agents’ current information $\mathbf{a}_{t}^{i}$ and map information $\mathbf{m}^{i}$, omitting agents’ past information. Second, dropout layers are added after each fully connected layer in the attention blocks and the MLP to enable BNN approximation using dropout [35].
@@ 4 Post-Optimizer
The epistemic uncertainty is inevitable for neural networks, including MotionFormer. MotionFormer may output dangerous actions in unfamiliar situations, e.g., a trajectory leading to a collision with a pedestrian. Therefore, a model-driven post-optimizer is developed to further optimize the action output by MotionFormer in the testing stage.
The post-optimizer consists of QP path planning and QP speed planning similar to Ref. [37], both of which are performed in the Frenet frame of the road. In the post-optimization, an action is decoupled into a path profile $y(x)$ and a speed profile $x(t)$, which are smoothing spline functions. Here, $x$, $y$, and $t$ represent the longitudinal coordinate, lateral coordinate, and time, respectively. Specifically, $y(x)$ is defined as
@@
where $y_{i}(x)=p_{i,0}+p_{i,1}x+\ldots p_{i,m}x^{m}$, $i=0,1,...,n-1$, is a $m$-th degree polynomial function with the coefficient $\boldsymbol{p}_{i}=\left(p_{i,0},p_{i,1},\ldots,p_{i,m}\right)^{\top}$, and $\{x_{0},x_{1},...,x_{n}\}$ are the spline knots. The parameter vector of the smoothing spline function is denoted as $\boldsymbol{p}=\left(\boldsymbol{p}_{0}^{\top},\boldsymbol{p}_{1}^{\top},\ldots,\boldsymbol{p}_{n-1}^{\top}\right)^{\top}$. $x(t)$ has the same form as $y(x)$.
@@
where $\hat{y}(x)$ represents the target path profile derived from the most likely trajectory output by MotionFormer. The linear constraints are related to collision avoidance, vehicle dynamics, and traffic regulations. For example, a constraint can be established based on the predicted trajectories of surrounding agents to avoid collisions.
@@ 5.1.2 Simulator and Metrics
To evaluate the performance of a planner, the closed-loop non-reactive score and the closed-loop reactive score were adopted following the standard evaluation protocol of the nuPlan dataset. Each score is a weighted average of metrics, including no at-fault collisions, drivable area compliance, making progress, driving direction compliance, time-to-collision within bound, progress along route ratio, speed limit compliance, and comfort.
@@ 5.2 Baselines
Urban Driver [40] and QCNet [27] were adopted as the baselines in our experiments. The Urban Driver model embeds vectorized state representation of agents and map polygons into local features which are then passed through a global attention layer to predict trajectories. The Urban Driver model was trained using open-loop imitation learning implemented by nuPlan.
QCNet, which won the championship of the Argoverse 2 multi-agent motion forecasting competition, processes query-centric state representation into scene encodings using a factorized attention mechanism and predicts multimodal trajectories with DETR-like decoders [41]. We implemented QCNet in the nuPlan framework and trained it using open-loop imitation learning. Furthermore, the size of QCNet was reduced to enable training on a single NVIDIA RTX 4090 GPU with 24GB of VRAM, where the Urban Driver model and our MotionFormer, CriticFormer, and Bayesian RewardFormer were also trained. In order to compensate for the performance loss caused by model size reduction, a masked training strategy [42] and a self-distillation strategy [43, 44] were used for training QCNet, resulting in a variant of QCNet namely query-centric masked autoencoder (QCMAE).
@@ 5.3 Implementation details
We first trained Bayesian RewardFormer using AVRL. Then, QCMAE was trained using the masked training and self-distillation strategies. Next, with the trained QCMAE as a warm-start, MotionFormer was trained using the proposed RITP method. The optimizer, learning rate scheduler, and hyperparameters used for training are summarised in Table 1. For more details, please refer to the source code at https://github.com/Zigned/nuplan_zigned.
@@
Batch size $N$ 1 4 4, $1$Footnote for Data 2
@@ 5.4 Comparison with Baselines
The results of the baselines and proposed approach on the testing set are summarized in Table 2. RITP involves training MotionFormer to predict the trajectories of all surrounding agents (referred to as multi-agent supervision hereafter), while RITP– focuses solely on optimizing the trajectory prediction for the ego agent. HRITP and HRITP–, where “H" stands for hybrid-driven, employ the post-optimizer to refine the output of MotionFormer.
@@
Closed-loop non-reactive Closed-loop reactive
[…]
Urban Driver[40] 0.53 0.48 0.53 0.32 0.56 0.50 0.58 0.29
QCMAE 0.54 0.49 0.55 0.36 0.63 0.57 0.66 0.38
RITP– 0.75 0.69 0.76 0.39 0.81 0.78 0.82 0.41
HRITP– 0.86 0.78 0.87 0.55 0.86 0.76 0.88 0.54
RITP 0.80 0.74 0.81 0.51 0.82 0.78 0.83 0.52
HRITP 0.89 0.81 0.90 0.62 0.89 0.82 0.90 0.61
[…]
  • NAFC: no at-fault collisions. TTCWB: time-to-collision within bound. SR: success rate (the number of scenarios without at-fault collisions divided by the number of all scenarios). $\uparrow$" denotes higher is better. The best result for each metric is in bold.
Comparing the results of Urban Driver, QCMAE, and HRITP in Table 2, it can be concluded that our HRITP method substantially outperforms the baselines in terms of both the closed-loop non-reactive score (over 72% improvement) and the closed-loop reactive score (over 60% improvement). Furthermore, HRITP significantly reduces at-fault collisions and better controls vehicle spacing, improving the success rate by over 63% and 36% for the two modes, respectively. Fewer collisions happened in the closed-loop reactive mode than in the closed-loop non-reactive mode because the surrounding agents in the latter mode automatically controlled the distance to the ego agent to avoid collisions.
The performance improvement of HRITP is attributed to several factors. Firstly, the superior performance of RITP over QCMAE demonstrates that RITP achieves better generalization through effective environment exploration based on the RL paradigm. Secondly, the comparison between RITP and HRITP highlights the effectiveness of the post-optimizer in refining MotionFormer’s output, thus preventing erroneous plans. Furthermore, the multi-agent supervision technique also contributes to performance improvement. This technique allows MotionFormer to fully utilize information from each scenario to model interactions between agents, thus improving the accuracy of trajectory prediction for the ego agent and surrounding agents. Note that predicted trajectories are crucial as they provide collision-avoidance constraints for the post-optimizer. We also found that integrating IL into RL is necessary. Without IL, MotionFormer trained solely by RL failed to generate a feasible trajectory and only produced a set of disorganized positions.

### v2

@@ Abstract
Reinforcement learning (RL) faces challenges in trajectory planning for urban automated driving due to the poor convergence of RL and the difficulty in designing reward functions. Consequently, few RL-based trajectory planning methods can achieve performance comparable to that of imitation learning-based methods. The convergence problem is alleviated by combining RL with supervised learning. However, most existing approaches only reason one step ahead and lack the capability to plan for multiple future steps. Besides, although inverse reinforcement learning holds promise for solving the reward function design issue, existing methods for automated driving impose a linear structure assumption on reward functions, making them difficult to apply to urban automated driving. In light of these challenges, this paper proposes a novel RL-based trajectory planning method that integrates RL with imitation learning to enable multi-step planning. Furthermore, a transformer-based Bayesian reward function is developed, providing effective reward signals for RL in urban scenarios. Moreover, a hybrid-driven trajectory planning framework is proposed to enhance safety and interpretability. The proposed methods were validated on the large-scale real-world urban automated driving nuPlan dataset. Evaluated using closed-loop metrics, the results demonstrated that the proposed method significantly outperformed the baseline employing the identical policy model structure and achieved competitive performance compared to the state-of-the-art method. The code is available at https://github.com/Zigned/nuplan_zigned.
@@ 1 Introduction
RL has been successfully applied in highway driving tasks, such as lane keeping [8, 9], lane changing [10, 11], and traffic merging [12], where the agents are usually responsible for steering, throttle, and braking control according to vectorized traffic information (e.g., relative positions and relative velocities). While RL has also been explored for urban scenarios, these environments involve more complex situations than highway scenarios, attributed to intricate road structures and diverse traffic participants with different behavioral patterns. This complexity necessitates more sophisticated state representations, leading to larger parameter spaces and more complex neural network architectures for RL agents. Consequently, the gradients computed by RL become insufficient to optimize such deep neural networks [13]. To address this, some approaches have combined RL with supervised learning. For instance, an IL agent was pre-trained to regularize the RL agent training by imposing the Kullback-Leibler divergence between the two agents’ policies in Ref. [14]. The RL agent determined a longitudinal target speed and a target lane according to bird’s-eye-view (BEV) rasterized images, which incorporate map information and relative states between traffic participants. Roach [15] trains an RL agent with privileged BEV semantic segmentation images. Then, the RL agent is employed to collect a dataset, with which a downstream IL agent learns to control the steering, throttle, and braking. More recently, GIGAFLOW [16] enables effective training of policies via self-play reinforcement learning on a massive scale. In GIGAFLOW, all traffic participants share the same policy and collecting massive experience in parallel, which mitigates the problem of insufficient gradient information.
However, few RL-based methods for AD are competitive to IL-based methods, especially in urban scenarios [13]. Beyond the aforementioned convergence challenges, another significant factor is that most existing RL methods for AD reason only one step ahead and lack the capability of planning for multiple future steps [7, 13, 17, 18], which can be necessary to handle complex urban scenarios. To enable multi-step reasoning, TrajHF [19] trains a diffusion model via reinforcement learning with human feedback to generate multiple future states. However, the iterative denoising process required for state generation can be computationally expensive during inference. As an alternative, trajectory value learning (TRAVL) [7] defines the action in RL as trajectories, which circumvents the need for explicit multi-step planning by directly learning the value of entire trajectories in RL. Nevertheless, TRAVL has several limitations. First, TRAVL selects a trajectory from a finite trajectory set generated by predefined rules, which may reduce trajectory optimality. Second, generating the trajectory set and evaluating every trajectory online increase computational complexity. Third, the applied scenarios of TRAVL were highway scenarios, while urban scenarios are much more complex.
@@ 2.1 Preliminaries
For high-dimensional problems, such as AD, the value function is usually approximated by a neural network $Q_{\theta}(s,a)$ parameterized by $\theta$. The deep Q-learning algorithm [26] optimizes the approximator $Q_{\theta}(s,a)$ to match a target $q$ outputted by a secondary frozen target network $Q_{\theta^{\prime}}(s,a)$:
@@ 2.2 Framework
As discussed in Section 1, training a deep neural network-based policy for urban trajectory planning using conventional RL methods is challenging, as the gradients provided by RL are insufficient for optimizing the policy. Given that IL offers better convergence properties, a natural solution is to combine RL with IL, using RL to explore the state and trajectory spaces, while employing IL to optimize the policy. However, IL requires ground truth, meaning that when the RL agent interacts with the environment, IL must access the optimal future trajectory for each state, which is intractable. This represents a key challenge of this approach. Therefore, we propose the RITP method, which integrates IL into RL within an actor-critic framework, as shown in Fig. 1. RITP incorporates an action sampler in addition to the actor to generate potential future trajectories, estimates their critic values using the critic network, and selects the optimal trajectory based on these values. The optimal trajectory is then used as a supervisory signal of IL to update the actor network.
Specifically, at each time step $t$, the actor takes an action $a_{t}^{0}$ given a state $s_{t}$. At the same time, the action sampler generates $S$ actions for the same state $s_{t}$, i.e., $\{a_{t}^{i}\}_{i=1}^{S}$, where $i$ is the index of the action. The actions are defined as trajectories $\{a_{t}^{i}\mid a_{t}^{i}=\{\mathbf{p}_{t+j}^{i}\}_{j=1}^{T_{\text{p}}}\}_{i=0}^{S}$, with spatial location $\mathbf{p}_{t+j}^{i}=(x^{i}_{t+j},y^{i}_{t+j})$ at $t+j$ time step, longitudinal and lateral coordinates $x^{i}_{t+j},y^{i}_{t+j}$, and planning time horizon $T_{\text{p}}$. Afterward, only the action generated by the actor, i.e., $a_{t}^{0}$, is sent to the environment, which then outputs the reward $r_{t+1}$ and the subsequent state $s_{t+1}$. Next, observing the state $s_{t+1}$, the actor and action sampler output actions $\{a_{t+1}^{i}\}_{i=0}^{S}$, and the process repeats accordingly.
@@
While the actor and the action sampler interact with the environment, the experience, defined as a tuple $(s_{t},\{a_{t}^{i}\}_{i=0}^{S},r_{t+1},s_{t+1})$, is added to the replay buffer at each time step. Subsequently, a mini-batch of experiences is sampled from the replay buffer to update the critic network. Afterward, for each experience, the action with the highest state-action value estimated by the critic network is selected from $\{a_{t}^{i}\}_{i=0}^{S}$ as the optimal action $a_{t}^{*}$. Finally, the optimal state-action pair $(s_{t},a_{t}^{*})$ is used to update the actor via IL, and the updated actor is employed to interact with the environment at the next time step.
@@ 2.3 Environment
The environment can be built on a real-world urban automated driving dataset. During training, the environment is in non-reactive log-replay mode, where the other traffic participants, such as vehicles, pedestrians, and cyclists, do not react to the behavior of the ego agent and simply act according to the logged data. Upon receiving the action $a_{t}^{0}$ from the actor, the environment propagates the motion of the ego vehicle using an LQR-based trajectory tracker [28] and the kinematic bicycle model. Note that an action is defined as a trajectory to be tracked by a downstream controller, rather than as multi-step control signals, primarily for the convenience of action sampling and post-processing. Additionally, the LQR-based trajectory tracker can be substituted with other controllers, such as a model predictive controller.
@@ 2.6 Action Sampler
Ideally, the actor should imitate the optimal action $a^{*}$ that maximizes the critic given a state $s$, i.e., $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$. However, finding the optimal action is our goal in the first place. It is intractable because the action space $\mathcal{A}$ for AD is continuous, high-dimensional, and with kinematic and dynamic constraints. Therefore, TRAVL [7] samples spline-based actions that are physically feasible for the ego agent in the Frenet frame of the road [31]. Inspired by TRAVL, we propose sampling actions based on piecewise polynomials, allowing more flexible driving maneuvers.
@@
Figure 4: Sampling actions based on piecewise polynomials
@@ 2.8 Training Objectives
In practice, the MotionFormer is optimized to predict not only the ego agent’s trajectory but also the trajectories of surrounding agents, whose supervisory signals are derived from the dataset. Because the predicted trajectories for surrounding agents are useful for the post-optimization (see Section 4). This training strategy is referred to as multi-agent supervision hereafter.
@@ 3.1 Model Structure
Bayesian RewardFormer, as shown in Fig. 16, has the same architecture as the actor network’s encoder except for two modifications. First, the temporal attention block is removed because Bayesian RewardFormer takes as input agents’ current information $\mathbf{a}_{t}^{i}$ and map information $\mathbf{m}^{i}$, omitting agents’ past information. Second, dropout layers are added before each fully connected layer in the attention blocks and the MLP to enable BNN approximation using dropout [37].
@@ 4 Post-Optimizer
The epistemic uncertainty is inevitable for neural networks, including MotionFormer. MotionFormer may output dangerous actions in unfamiliar situations, e.g., trajectories leading to pedestrian collisions. Therefore, a model-driven post-optimizer is developed to further refine the action outputted by MotionFormer in the testing stage.
The post-optimizer comprises trajectory proposal generation, QP path planning, and QP speed planning, all performed within the road’s Frenet frame. In the trajectory proposal generation step, the post-optimizer processes the action $a$ outputted by MotionFormer into a trajectory proposal. Specifically, candidate trajectories (up to 1,500 trajectories in our experiments) are generated based on piecewise polynomials as shown in Fig. 4. Subsequently, a closed-loop simulation is conducted for these candidate trajectories using an LQR trajectory tracker and a kinematic bicycle model to obtain simulated rollouts, following the method described in Ref. [39]. Each rollout is then evaluated based on criteria such as driving safety, traffic rule compliance, and progress along route, utilizing a rule-based evaluation framework established in Ref. [39]. This evaluation assigns each candidate trajectory a rule-based score. The final score for each candidate trajectory is calculated by adding the rule-based score and a preference score, which is defined as the L2 distance between the action $a$ and the candidate trajectory. Finally, the trajectory proposal is selected as the candidate trajectory with the highest final score. The motivation for this step is to preliminarily refine potentially dangerous actions outputted by MotionFormer, which could otherwise mislead the subsequent QP path and speed planning.
However, the generated trajectory proposals may be inflexible in handling diverse situations due to their predefined shapes. Therefore, QP path and speed planning are necessary to further refine the trajectory proposal. Specifically, a trajectory proposal is decoupled into a path profile $y(x)$ and a speed profile $x(t)$, represented as smoothing spline functions. Here, $x$, $y$, and $t$ represent the longitudinal coordinate, lateral coordinate, and time, respectively. Specifically, $y(x)$ is defined as
@@
where $y_{i}(x)=p_{i,0}+p_{i,1}x+\ldots+p_{i,m}x^{m}$, $i=0,1,...,n-1$, is a $m$-th degree polynomial function with the coefficient $\boldsymbol{p}_{i}=\left(p_{i,0},p_{i,1},\ldots,p_{i,m}\right)^{\top}$, and $\{x_{0},x_{1},...,x_{n}\}$ are the spline knots. The parameter vector of the smoothing spline function is denoted as $\boldsymbol{p}=\left(\boldsymbol{p}_{0}^{\top},\boldsymbol{p}_{1}^{\top},\ldots,\boldsymbol{p}_{n-1}^{\top}\right)^{\top}$. $x(t)$ has the same form as $y(x)$.
@@
where $\hat{y}(x)$ represents the target path profile derived from the most likely trajectory outputted by MotionFormer. The linear constraints are related to collision avoidance, vehicle dynamics, and traffic regulations. For example, a constraint can be established based on the predicted trajectories of surrounding agents to avoid collisions.
@@ 5.1.2 Simulator and Metrics
To evaluate the performance of a planner, the closed-loop non-reactive score and the closed-loop reactive score were adopted following the standard evaluation protocol of the nuPlan dataset. Each score is a weighted average of metrics, including no at-fault collisions, drivable area compliance, making progress, driving direction compliance, time-to-collision within bound, progress along route ratio, speed limit compliance, and comfort. Additionally, we defined a metric of success rate as the number of scenarios without at-fault collisions divided by the total number of scenarios.
@@ 5.2 Baselines
Urban Driver [41], QCNet [29], and PLUTO [42] were adopted as the baselines in our experiments. The Urban Driver model embeds vectorized state representation of agents and map polygons into local features which are then passed through a global attention layer to predict trajectories. The Urban Driver model was trained using open-loop imitation learning implemented by nuPlan.
QCNet [29], which won the championship of the Argoverse 2 multi-agent motion forecasting competition, processes query-centric state representation into scene encodings using a factorized attention mechanism and predicts multimodal trajectories with DETR-like decoders [43]. We implemented QCNet in the nuPlan framework and trained it using open-loop imitation learning. Furthermore, the size of QCNet was reduced to enable training on a single NVIDIA RTX 4090 GPU with 24GB of VRAM, where the Urban Driver model and our MotionFormer, CriticFormer, and Bayesian RewardFormer were also trained. In order to compensate for the performance loss caused by model size reduction, a masked training strategy [44] and a self-distillation strategy [45, 46] were used for training QCNet, resulting in a variant of QCNet namely query-centric masked autoencoder (QCMAE).
PLUTO [42] is a state-of-the-art (SOTA) planning method on the nuPlan dataset. It uses a longitudinal-lateral aware transformer with factorized self-attention to jointly model maneuvers and generate diverse, reference-line-conditioned trajectory proposals for flexible and safe driving. Furthermore, PLUTO employs contrastive imitation learning with novel data augmentations (e.g., state perturbation, agent dropout) to mitigate causal confusion and distribution shift commonly faced by traditional IL methods. Additionally, PLUTO integrates a rule-based safety verification module to improve driving safety. We trained PLUTO using the hyperparameters specified in Ref. [42], on the same dataset and hardware setup as our RITP model and other baseline methods.
@@ 5.3 Implementation details
We first trained Bayesian RewardFormer using AVRL. Then, QCMAE was trained using the masked training and self-distillation strategies. Next, with the trained QCMAE as a warm start, MotionFormer was trained using the proposed RITP method. The optimizer, learning rate scheduler, and hyperparameters used for training are summarised in Table 1. For more details, please refer to the source code at https://github.com/Zigned/nuplan_zigned.
@@
Batch size $N$ 1 4 4, 11
@@ 5.4 Comparison with Baselines
The results of the baselines and proposed approach on the testing set are summarized in Table 2. Comparing the results, it can be concluded that our RITP method substantially outperforms the Urban Driver and QCMAE in terms of both the closed-loop non-reactive score (over 125.0% improvement) and the closed-loop reactive score (over 113.1% improvement). This remarkable improvement over QCMAE is particularly noteworthy given that RITP shares the same policy network structure, highlighting the effectiveness of the proposed method. Furthermore, RITP achieved competitive composite scores compared to that of the SOTA method PLUTO, even outperforming PLUTO on safety-related metrics (NAFC, TTCWB, and, SR). The slight decrease in scores on metrics related to traffic rules and driving tasks (DAC, DDC, and PARR) is deemed acceptable given the paramount importance of safety in automated driving. Regarding the comfort metric CMFT, RITP and PLUTO are both inferior to UrbanDriver and QCMAE. This is a consequence of compromises made in their design to handle challenging emergency situations (e.g., hard braking). Note that fewer collisions happened in the closed-loop reactive mode than in the closed-loop non-reactive mode because the surrounding agents in the latter mode automatically controlled the distance to the ego agent to avoid collisions.
@@
Method Mode NAFC TTCWB DAC DDC SLC MP PARR CMFT SR score
Urban Driver[41] CL-NR 0.53 0.48 0.50 0.69 0.92 0.95 0.80 0.97 0.53 0.32
QCMAE 0.54 0.49 0.68 0.95 0.98 0.96 0.76 0.99 0.55 0.36
RITP (ours) 0.94 0.90 0.97 0.99 1.00 0.96 0.83 0.81 0.94 0.81
PLUTO [42] 0.93 0.86 0.99 0.99 0.98 0.98 0.91 0.79 0.93 0.84
Urban Driver[41] CL-R 0.56 0.50 0.51 0.69 0.92 0.96 0.82 0.97 0.58 0.29
QCMAE 0.63 0.57 0.67 0.94 0.98 0.96 0.75 0.99 0.66 0.38
RITP (ours) 0.95 0.89 0.97 0.99 1.00 0.98 0.80 0.82 0.95 0.81
PLUTO [42] 0.94 0.88 1.00 0.99 0.98 0.97 0.85 0.77 0.95 0.83
  • CL-NR: closed-loop non-reactive. CL-R: closed-loop reactive. NAFC: no at-fault collisions. TTCWB: time-to-collision within bound. DAC: drivable area compliance. DDC: driving direction compliance. SLC: speed limit compliance. MP: making progress. PARR: progress along route ratio. CMFT: comfort. SR: success rate. Higher is better for all metrics. The best result for each metric is in bold, and the second best result is underlined.